# AQORA: A Learned Adaptive Query Optimizer for Spark SQL

### v1

@@ Abstract
Recent studies have identified two main approaches to improve query optimization: learned query optimization (LQO), which generates or selects better query plans before execution based on models trained in advance, and adaptive query processing (AQP), which adapts the query plan during execution based on statistical feedback collected at runtime. Although both approaches have shown promise, they also face critical limitations. LQO must commit to a fixed plan without access to actual cardinalities and typically rely on a single end-to-end feedback signal, making learning inefficient. On the other hand, AQP depends heavily on rule-based heuristics and lacks the ability to learn from experience. In this paper, we present AQORA, an adaptive query optimizer with a reinforcement learning architecture that combines the strengths of both LQO and AQP. AQORA addresses the above challenges through four core strategies: (1) realistic feature encoding, (2) query stage-level feedback and intervention, (3) automatic strategy adaptation, and (4) low-cost integration. Experiments show that AQORA reduces end-to-end execution time by up to 90% compared to other learned methods and by up to 70% compared to Spark SQL’s default configuration with adaptive query execution.
@@ I Introduction
This leads to a fundamental question: can we design a system that effectively combines the strengths of pre-execution learning with in-execution adaptation? Fig. 1 illustrates the design space of query optimizers along two key axes: when optimization occurs (pre-execution vs. in-execution) and how decisions are made (rule-based vs. learning-based). Most LQOs, such as AutoSteer [9] and Lero [10], perform the optimization in the pre-execution phase and make decisions without access to runtime feedback. In contrast, AQP methods, such as Spark SQL’s Adaptive Query Execution and runtime filters, perform limited rule-based adjustments during execution. As shown in the figure, our proposed system, AQORA (Adaptive Query Optimizer with Reinforcement learning Architecture), is the first to reside in the top-right quadrant: a learned query optimizer that performs in-execution plan adaption on Spark SQL. It integrates with Spark’s runtime engine via a lightweight planner extension and uses an actor-critic reinforcement learning model to make informed decisions during execution, leveraging query stage-level feedback for fine-grained supervision. It is designed as a plug-and-play module, requiring minimal changes to the underlying system, and can be easily integrated into Spark SQL.
@@ II Related Work
In this section, we review related work, including LQOs and previous AQP methods. Since the pioneering work of LEO [29] in 2001, which introduced the concept of a learning-based query optimizer, recent studies on LQO have evolved into two main directions: plan-constructors and plan-steerers.
@@ IV System Overview
As shown in Fig. 5, AQORA consists of two core components: a planner extension and a decision model. The workflow proceeds as follows:
@@
    When Spark SQL’s AQE is triggered during execution, the planner extension extracts the current query plan along with actual cardinality statistics and forwards it to the decision model. This process can also be triggered during the query optimization phase, in which case actual cardinalities are not available.
@@
    The decision model analyzes the input plan and returns optimization actions (e.g., swapping the join order of leaf nodes).
@@
    The planner applies these actions to the ongoing query plan and resumes execution using the optimized version.
AQORA iteratively performs steps 1)-3) until the query completes and then performs step 4). We formalize this $k$-step process as a trajectory:
@@
4) After the query completes, the decision model replays the collected trajectory to learn from past decisions and improve its optimization strategy for future queries.
@@ V-A1 State, Action, and Return
The action space of AQORA is carefully designed to balance flexibility and tractability in adaptive query optimization. The first type (CBO action) is included because the quality of the initial query plan significantly impacts the subsequent optimization trajectory. Dynamically controlling this setting allows the agent to make better trade-offs between optimization quality and performance cost.
Once an initial plan is chosen, AQORA can refine it through two kinds of structural modifications: adjusting the join order and modifying the join method. The lead action provides a more practical and efficient mechanism for guiding join order, as most beneficial changes can be achieved by repositioning a single leaf node to the front.
For join method selection, the key issue is the overly conservative application of BJT. To address this, AQORA provides the broadcast action, which forces a broadcast join on a selected leaf node. While this strategy introduces some risk, it enables effective correction of conservative join method choices.
Finally, the no-op action allows the model to intentionally defer decision and wait for more runtime information.
@@
where $T_{\text{execute}}$ represents the total execution time of $\tau$ (in seconds). The square root operation helps mitigate reward scale variations across different queries. If query execution fails, $R(\tau)$ is assigned a substantial negative penalty (e.g., $\sqrt{300}$). The intermediate reward $r_{i}$ is defined as the negative value of the additional number of shuffles introduced after taking action $a_{i}$, normalized by 10. This definition naturally provides positive feedback when the model selects no-op (no additional shuffle is introduced), encouraging safe behavior during early training, while penalizing actions that introduce extra shuffles and discouraging costly changes late in the optimization process. The discount rate $\gamma(<1)$ ensures that the accumulated intermediate rewards $r_{i}$, even over many steps, do not dominate the overall cost $\sqrt{T_{\text{execute}}}$.
@@ V-B2 Feature Encoding
  • Each node is assigned a type, such as a join or a broadcast query stage. AQORA encodes the node type as a one-hot vector type(u)(u). In a simplified example shown in Figure 6, if $u$ is a leaf node, then type(u)=(0,1)(u)=(0,1); if $u$ is a join node, then type(u)=(1,0)(u)=(1,0)
@@
The encoded vector tree is processed by AQORA’s actor and critic networks.
We choose TreeCNN for its low optimization overhead and superior performance after being trained on a sufficient number of queries (2400 on ExtJOB).
@@ VII-D4 Different Action Space
Although AQORA’s planner extension supports a wide range of action types, we conducted ablation studies to identify which subsets of the action space contribute most to learning efficiency and convergence. Specifically, we experimented with enabling or disabling the broadcast hint, the leading hint, and the CBO toggle. As shown in the results, enabling the broadcast hint often leads the model to generate plans that broadcast excessively large tables during training, resulting in significant performance fluctuations and instability.
When leading actions are disabled, AQORA loses its ability to explicitly manipulate join ordering, thereby reducing its optimization power and leading to suboptimal plans for many complex queries. Finally, disabling the CBO action prevents AQORA from jumping to an entirely different planning path starting from a CBO-generated plan. This restriction increases the difficulty for AQORA to escape local optima and discover globally better execution plans.

### v2

# AQORA: A Fast Learned Adaptive Query Optimizer with Stage-Level Feedback for Spark SQL

@@ Abstract
Recent studies have identified two main approaches to improve query optimization: learned query optimization (LQO), which generates or selects better query plans before execution based on models trained in advance, and adaptive query processing (AQP), which adapts the query plan during execution based on statistical feedback collected at runtime. Although both approaches have shown promise, they also face critical limitations. LQO must commit to a fixed plan without access to actual cardinalities and typically rely on a single end-to-end feedback signal, making learning inefficient. On the other hand, AQP depends heavily on rule-based heuristics and lacks the ability to learn from experience. In this paper, we present AQORA, a learned adaptive query optimizer that continuously refines query plans through stage-level feedback, unifying pre-execution and in-execution optimization within a single learning process. AQORA addresses the above challenges through four core strategies: (1) realistic feature encoding, (2) query stage-level feedback and intervention, (3) automatic strategy adaptation, and (4) low-cost integration. Experiments show that AQORA reduces end-to-end execution time by up to 90% compared to other learned methods and by up to 70% compared to Spark SQL’s default configuration with adaptive query execution.
@@ I Introduction
This leads to a fundamental question: can we design a query optimizer that continuously learns and adapts across both the pre- and in-execution phases? Fig. 1 illustrates the design space of query optimizers along two key axes: when optimization occurs (pre-execution vs. in-execution) and how decisions are made (rule-based vs. learning-based). Most LQOs, such as AutoSteer [9] and Lero [10], perform the optimization in the pre-execution phase and make decisions without access to runtime feedback. In contrast, AQP methods, such as Spark SQL’s adaptive query execution and runtime filters, perform limited rule-based adjustments during execution. As shown in the figure, our proposed system, AQORA (Adaptive Query Optimizer with Reinforcement learning Architecture), is the first to reside in the top-right quadrant: a learned adaptive query optimizer that performs in-execution plan refinement on Spark SQL. It integrates with Spark’s runtime engine via a lightweight planner extension and uses an actor-critic reinforcement learning model to make informed decisions during execution, leveraging query stage-level feedback for fine-grained supervision. It is designed as a plug-and-play module, requiring minimal changes to the underlying system, and can be easily integrated into Spark SQL.
@@ II Related Work
In this section, we review related work on LQO and AQP. Since the pioneering work of LEO [29] in 2001, which introduced the concept of a learning-based query optimizer, recent studies on LQO have evolved into two main directions: plan-constructors and plan-steerers.
@@ IV System Overview
As shown in Fig. 5, AQORA consists of two core components: a planner extension and a decision model. The figure illustrates an example where AQORA performs an optimization step during the execution of the STACK query q1#1.
@@
    When Spark SQL’s AQE is triggered during execution, the planner extension extracts the current query plan along with actual cardinality statistics and forwards it to the decision model. In Fig. 5, the system has just completed scanning table t and observed that the output cardinality is only 1. The planner extension then sends the partially executed plan and the row count (1) to the decision model. This process can also be triggered during the query optimization phase, in which case actual cardinalities are not available.
@@
    The decision model analyzes the input plan and returns optimization actions. As shown in Fig. 5, the decision model observes that table t produces only one output row but is joined after two larger tables. The actor therefore infers that placing t earlier in the join order could be more efficient and outputs the action lead(t).
@@
[…]
4) After the query completes after $k$ iterations of 1)-3), the decision model replays the collected $k$ step optimization process to learn from past decisions and improve its optimization strategy for future queries.
We formalize this optimization process as a trajectory:
@@ V-A1 State, Action, and Return
The action space of AQORA is carefully designed to balance flexibility and tractability in adaptive query optimization. In some cases, enabling CBO produces a better starting plan that is closer to the optimal; in others, disabling it yields better results. Once an initial plan is chosen, AQORA can refine it through two kinds of structural modifications: adjusting the join order and modifying the join method. The lead action provides a more practical and efficient mechanism for guiding join order, as most beneficial changes can be achieved by repositioning a single leaf node to the front. Finally, the no-op action allows the model to intentionally defer decision and wait for more runtime information.
@@
where $T_{\text{execute}}$ represents the total execution time of $\tau$ (in seconds). The square root operation helps mitigate reward scale variations across different queries. If query execution fails, $R(\tau)$ is assigned a substantial negative penalty (e.g., $\sqrt{300}$). The intermediate reward $r_{i}$ is defined as the negative value of the additional number of shuffles introduced after taking action $a_{i}$, normalized by 10. This definition naturally provides positive feedback when the model selects no-op (no additional shuffle is introduced), encouraging safe behavior during early training, while penalizing actions that introduce extra shuffles and discouraging costly changes late in the optimization process. The discount rate $\gamma(<1)$ ensures that the accumulated intermediate rewards $r_{i}$, even over many steps, do not dominate the overall cost $R(\tau)$.
@@ V-B2 Feature Encoding
  • Each node is assigned a type, such as a join or a broadcast query stage. AQORA encodes the node type as a one-hot vector type(u)(u). In a simplified example shown in Fig. 6, if $u$ is a leaf node, then type(u)=(0,1)(u)=(0,1); if $u$ is a join node, then type(u)=(1,0)(u)=(1,0)
@@
The encoded vector tree is then processed by AQORA’s actor and critic networks, and we choose TreeCNN for its low optimization overhead and superior performance after being trained on a sufficient number of queries (2400 on ExtJOB).
@@ VII-D4 Different Action Space
AQORA’s planner extension supports a wide range of action types, we conducted ablation studies to identify which subsets of the action space contribute most to learning efficiency and convergence. As shown in the results, enabling the broadcast hint often leads the model to generate plans that broadcast excessively large tables during training, resulting in significant performance fluctuations and instability. When leading actions are disabled, AQORA loses its ability to explicitly manipulate join ordering, thereby reducing its optimization power and leading to suboptimal plans for many complex queries. Finally, disabling the CBO action prevents AQORA from jumping to an entirely different planning path starting from a CBO-generated plan. This restriction increases the difficulty for AQORA to escape local optima and discover globally better execution plans.
@@ VIII Conclusion
IX AI-Generated Content Acknowledgement
Portions of this paper’s text were refined using ChatGPT 5 to improve the clarity, grammar, and conciseness of the language. The AI system was used only for linguistic polishing and did not contribute to the conceptual development, technical design, experimental results presented in this work.